|
ID |
Date |
Author |
Subject |
Text |
 |
|
479
|
Mon Jan 29 08:52:07 2024 |
TD JM | Monday 29 January |
09.50 CARME IE514 5.1e-11 mbar ( ambient
temperature 16 deg C )
|
32x |
|
478
|
Sat Jan 27 20:38:15 2024 |
PJCS | [INFO] Steps to diagnose White Rabbit problems | This is a suggested sequence of actions
to help with the current problem with the
White Rabbit timestamp distribution. Test |
|
|
477
|
Mon Jan 22 16:29:31 2024 |
TD | Saturday 20 January | 11.00 No WLAN/WiFi
came-gsi private LAN appears to be
OK
|
30x |
|
476
|
Fri Jan 19 11:17:50 2024 |
JG, TD | Friday 19 January |
12.20 CARME IE514 4.1e-11mbar, ambient temperature
16.7 deg C
|
18x |
|
475
|
Thu Jan 18 17:21:00 2024 |
TD | Thursday 18 January | Network ( dhcpd.conf ) and MIDAS startup.tcl/Options/ASIC
settings/rfs filesystems configured
|
|
|
474
|
Wed Jan 17 19:21:20 2024 |
TD | AIDA spares |
Spare AIDA hardware
|
|
|
473
|
Wed Jan 17 19:15:47 2024 |
TD | Wednesday 17 January |
CAEN N1419ET remote control via carme-gsi
and nnrpi1 - see https://elog.ph.ed.ac.uk/CARME/472
|
|
|
472
|
Wed Jan 17 10:31:05 2024 |
TD | Remote control of CAEN N1419ET from nnrpi1 |
Login to nnrpi1 from carme-gsi using ssh
|
|
|
471
|
Tue Jan 16 16:01:48 2024 |
TD | Change CARMEraspberrypi to nnrpi1 |
Change RPi from CARMEraspberrypi to nnrpi1
to provide carme-gsi access to Rly16 to control |
|
|
470
|
Tue Jan 16 14:36:30 2024 |
TD | CARME configuration |
FEE64
|
    |
|
469
|
Tue Jan 16 08:57:44 2024 |
JG, DR, TD | Tuesday 16 January | 10.00 Outflow manifold returned from workshop
with leaking attachment point welding re-worked.
|
8x |
|
468
|
Mon Jan 15 18:51:28 2024 |
JG, DR, TD | Monday 15 January |
Cooling water manifolds
|
|
|
467
|
Sun Jan 14 12:31:32 2024 |
JM, CB | FEE positions mac addresses | *Updated following FEE changes. Correct
as of 31/01/24
FEE modules are labelled according |
|
|
466
|
Sat Jan 13 14:32:21 2024 |
CB, JM | Leaky manifold and cable management | Found two leaks in the welding of the manifolds,
marked them in tape. Manifold to be repaired
on Monday. Remember to re-install missing |
  |
|
465
|
Sat Jan 13 14:25:05 2024 |
CB, JM | Bias tests with all FEEs connected | Connected all FEEs.
The two top fees corresponding
to the bottom right detector are connected |
 |
|
464
|
Thu Jan 11 19:57:20 2024 |
CB, JM | All FEE holders mounted | After discussions with Davide, changed
the holes in the FEE holders for the bottom-right
detector to slots in order to slightly change |
|
|
463
|
Wed Jan 10 19:27:08 2024 |
CB, JM | Manifold and more bias tests | Mounted water manifold on CARME with help
from GSI plumber. Connected all the hoses
with Swagelok self-locking connectors. Tested |
|
|
462
|
Wed Jan 10 10:22:41 2024 |
CB, JM | Some FEEs mounted - issues with Det#2 (bottom left) | Started mounting some FEEs. We ran into
some mechanical issues due to the new ERNI
connectors on the new flexi-rigid adaptor |
 |
|
461
|
Mon Jan 8 19:29:57 2024 |
CB, JM | Manifold repaired + fee assembly in bench | Davide extracted the stuck blind plugs
and re-threaded all missing ports. Completed
manifold assembly, except for main inlet/outlet |
   |
|
460
|
Mon Jan 8 19:27:36 2024 |
CB, JM | Leak current tests overnight | Left bias in overnight. Leakage current
decreased and stabilised. See attached. Bias
off as we prepare to mount the fees. |
|